# DeepProofLog: Efficient Proving in Deep Stochastic Logic Programs

### v1

@@ 5 Learning in DeepProofLog
Learning in DeepProofLog is equivalent to learning the optimal policy of the corresponding MDP defined in Section 4. Let the objective be $J(\pi_{\lambda})=\sum_{i=1}^{N}[V^{\pi_{\lambda}}(q^{(i)},y^{(i)})]$ (sutton1999reinforcement). Given our reward design in Equation 3 and setting the discount factor $\gamma=1$, we show in Appendix A.3 that the objective function simplifies to:
@@
Approximate Inference via Policy Gradient. When the full goal space is intractable, the mapping between SLP derivations and MDPs enables the application of reinforcement learning techniques to learn the clause selection policy. Prominent algorithms such as REINFORCE (williams1992simple), Actor-Critic (sutton2000policy), and Proximal Policy Optimization (PPO) (schulman2017proximal) can be adapted. For example, we employ PPO to optimize the neural policy, using the standard clipped surrogate objective. Alongside the policy, we parameterize a neural value function with the same architecture to estimate returns and reduce variance during training. While value-based methods are also applicable in principle, we leave their exploration to future work.
@@ 6 Experiments
We present two sets of experiments. First, we assess our approach on a neurosymbolic task that learns the perception given the background knowledge, showing that DeepProofLog (DPrL) outperforms advanced approximate inference systems in both accuracy and scalability, while scaling better than exact inference systems like DeepStochLog. The second set of experiments addresses explainable knowledge graph (KG) completion. Here, DPrL learns from data to generate high-probability proofs for positive queries and low-probability proofs for negative ones. The fact that each classification is expressed by a proof makes DPrL decisions fully interpretable. Comparative experiments against state-of-the-art proof-based KG systems show that our method scales to large KGs without compromising its predictive performance. All code and data will be publicly available upon publication.
@@ 6.1 MNIST Addition
NeSy Exact
KLay Time T/O
@@
NeSy Approximate
Scallop Acc. $92.3\pm 0.7$ T/O
@@
We compare with state-of-the-art NeSy methods along two dimensions: predicted sum accuracy and training time. The baselines include three exact systems: KLay (DBLP:conf/iclr/MaeneDM25), DeepStochLog, and DeepSoftLog (maene2023soft); and three approximate systems: Scallop (huang2021scallop), A-NeSI (van2023nesi), and EXAL (verreet2024explain). The results are summarized in Table 1, with details on the experimental setups and hyperparameters provided in Appendix C.1.
The DP variant achieves accuracy comparable to existing exact NeSy systems, while significantly outperforming both exact and approximate baselines in terms of training efficiency. We observed that no other previous method successfully scales to $N=100$ without timing out. In the approximate setting, our PG variant achieves the highest accuracy. It scales beyond $N=100$, demonstrating strong computational efficiency, but converges to a suboptimal policy due to the vast combinatorial search space.
6.2 Explainable Knowledge Graph Completion
@@
[…]
synset_domain_topic_of
(09788237, 08441203)
hypernym
(09788237, 09774266)
synset_domain_topic_of
(09774266, 08441203)
derivationally_related_form
(09774266, 00872886)
synset_domain_topic_of
(00872886, 08441203)
derivationally_related_form
(00872886, 09774266)
synset_domain_topic_of
(09774266, 08441203) $\square$
[…]
uncle(2266, 2252)
brother(2260, 2252)
uncle(2266, 2260)
sister(2262, 2260)
uncle(2266, 2262) $\square$
@@
Table 3: Examples of multi-step proof trees discovered by DeepProofLog for the Family and WN18RR datasets.
@@
We evaluate our approach on popular benchmarks for KG completion: Family (kok2007statistical) and WN18RR (dettmers2018convolutional). The Family dataset focuses on modeling familial relationships, and the WN18RR dataset is a subset of WordNet (WN18) created to provide a more challenging benchmark by removing inverse relations. More information about the dataset statistics can be found in the Appendix C.2. All methods have been tested with 200 randomly sampled negative corruptions per query.
@@
Table 2 compares DPrL against both KGE baselines (ComplEx (trouillon2016complex) and RotatE (sun2019rotate)) and other logic-based NeSy systems like Semantic-based Regularization (SBR) (diligenti2017semantic), R2N (marra2025relational_reasoning_networks) and DCR (barbiero2023interpretable). Please note that these NeSy systems would not scale to the considered datasets without relying on custom approximate grounding techniques (groundingmethods), which make it possible to perform approximate reasoning using these methodologies. We indicate these variants as $\textit{R2N}_{\text{SG}}$, $\textit{DCR}_{\text{SG}}$, and $\textit{SBR}_{\text{SG}}$. However, these fast grounding techniques require a manual fine-tuning of their parameters, whereas DPrL automatically learns the best exploration strategy for each dataset.
@@
DPrL uses Proximal Policy Optimization (PPO) (schulman2017proximal) to navigate the complex, large-scale action spaces emerging in KG reasoning. The model computes the scores $p^{\lambda}_{\text{success}}(q)$ for each query $q$ so that the standard evaluation metrics for each task (MRR, Hits@N) can be computed (see Appendix C.2 for the metric definitions). All NeSy methods, including DPrL, use RotatE output as a prior, learning a logic-based adjustment on top of it.
@@
For the Family dataset, all NeSy systems improve performance over the baseline. DPrL is particularly effective, especially on MRR and Hits@1 metrics. A similar trend is observed in the WN18RR dataset, where DPrL and SBR outperformed other methods.
@@
Unlike the partially latent reasoning done by R2N or DCR, DPrL is fully interpretable as each classification decision is the result of the execution of a proof, which can be inspected by a human. Table 3 showcases two examples of explanations in the form of proofs discovered by DPrL: a depth-3 proof for the query synset_domain_topic_of(09788237, 08441203) from the WN18RR dataset, and a proof for the query uncle(2266, 2252) in the Family dataset.
@@
Methods based on a complete or approximate grounding technique, like the ones proposed by (groundingmethods), would need to instantiate and process a combinatorially large number of ground formulas to find a proof at high depth. In contrast, DPrL can avoid an exhaustive search by learning an efficient proof-finding policy, which efficiently navigates deep reasoning paths. For example, on the Family dataset, the smart grounding techniques proposed by (groundingmethods) instantiate 18071, 45466, 181095 groundings for the test queries, when allowing proofs at depth 1,2,3, respectively. On the other hand, DPrL needs to evaluate only 12334 ground formulas to perform full inference over the test set.
@@
In conclusion, DPrL demonstrates superior scalability over other NeSy models capable of learning perception and structure, such as DeepSoftLog (which fails to scale to datasets presented here), while preserving the explainability and interpretability advantages of logic-based NeSy approaches.
@@ 7 Related Work
DeepStochLog extends SLPs by using neural networks to parameterize clause probabilities under derivation-based semantics. However, its exact inference leads to the combinatorial explosion of proof paths, making it intractable for complex tasks. It also assumes fixed output domains and conditions neural predictions on partial goals, limiting flexibility and context awareness. In contrast, DeepProofLog factorizes the derivation distribution over resolution steps, conditions each decision on the full current goal, and supports variable output domains. This results in better scalability (linear in derivation steps and matching unification), more informed decisions, and greater expressivity.
@@
Efforts to improve the scalability of exact NeSy systems include DeepStochLog, which adopts a more efficient derivation-based semantics, and KLay (DBLP:conf/iclr/MaeneDM25), which introduces GPU-parallelizable arithmetic circuits. Approximate NeSy methods include search-based techniques (manhaeve2021approximate; huang2021scallop) and random walks with restart (wang2013programming), which limit inference to parts of the proof space but risk bias. Sampling-based approaches (verreet2023explain) avoid full model counting but are still limited by grounding. Variational inference (abboud2020learning; dos2021neural; van2023nesi) replaces logic inference with learned predictors, trading interpretability for speed. Parametrized grounding (castellano2025grounding) offers a tunable balance between scalability and expressivity but lacks general guarantees like DeepProofLog.
@@
Automated Theorem Proving (ATP) aims to find a proof for a given query within deterministic logical frameworks. Recent advances in ATP integrate learning-based heuristics to guide proof search (rocktaschel2017end; kaliszyk2018reinforcement; rawson2019neurally; zombori2020prolog; lample2022hypertree). Unlike ATP systems, DeepProofLog operates in a probabilistic setting and learns from labeled data. Rather than determining provability, it optimizes inference by maximizing the probability of positive queries and minimizing that of negatives.
@@
Reinforcement learning (RL) has been applied to knowledge graph (KG) completion mainly via path-finding models like DeepPath (xiong2017deeppath) and MINERVA (das2017go), which train agents to navigate KGs by sequentially selecting entities and relations. These approaches map relation-path discovery to an MDP but are specific to KG structures. While paths can serve as explanations, they may be complex and hard to interpret. In contrast, DeepProofLog maps SLPs to MDPs, enabling RL-guided proof search across any domain expressible with SLPs, offering greater generality beyond KGs.
@@ 8 Conclusions
We introduce DeepProofLog, a scalable and flexible NeSy framework that bridges the gap between expressive logical representations and efficient, data-driven learning. DeepProofLog leverages SLPs with derivation-based semantics and introduces a more expressive neural parameterization by conditioning on the full goal at each resolution step. A primary contribution is the novel and formal correspondence between inference and learning in our deep SLPs and policy optimization in MDPs, enabling the use of dynamic programming and reinforcement learning techniques to tackle previously intractable NeSy problems.
The presented method shares a limitation with prior work: clause selection is restricted to the leftmost atom. While this preserves expressivity and completeness of the proofs, we plan to explore a more flexible selection, which could more efficiently prune failing derivations. Our current implementation uses standard policy gradients; future work will consider advanced RL techniques and broader evaluation settings. We also see potential in extending to possible world–based NeSy methods like Markov Logic Networks (richardson2006markov) and in studying automatic logic program reformulations to reduce the search space (leuschel2002logic).
9 Acknowledgments
This work has been supported by the EU Framework Program for Research and Innovation Horizon under the Grant Agreement No 101073307 (MSCA-DN LeMuR). F. Giannini acknowledges support from the Partnership Extended PE00000013 - FAIR - Future Artificial Intelligence Research - Spoke 1 Human-centered AI and ERC-2018-ADG G.A. 834756 XAI: Science and technology for the eXplanation of AI decision making. L. De Raedt is also supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Y. Jiao wants to thank Gabriele Venturato and Lennert De Smet for stimulating discussions.
@@ A.3 Alignment between MDP Objective and Cross-entropy Loss
In NeSy systems, learning is often supervised via cross-entropy between predicted success probabilities and binary labels (winters2022deepstochlog). We now show that, under our MDP formulation in Section 4, the objective that maximizes the expected cumulative reward is equivalent to minimizing cross-entropy loss on a labeled dataset $\mathcal{D}=\{(q^{(i)},y^{(i)})\}_{i=1}^{N}$.
@@ C.1 MNIST Addition
We follow the same data generation process for the MNIST addition experiment as in previous works (manhaeve2018deepproblog). This involves consecutively selecting $2N$ images from the MNIST dataset and splitting them into two distinct sequences, each of length $N$. To create supervision signals, we compute the target sum values by multiplying the labels of the selected MNIST images by the corresponding powers of 10 and summing the resulting sequences. Finally, the two numbers are added together to form the final label. Each MNIST image is used only once in all sequences, resulting in $\lfloor 60000/2N\rfloor$ learning samples. They are further split in a 5:1 ratio for training and validation. The test set is generated similarly using the MNIST test partition.
@@
As in previous works, we use a standard LeNet architecture in all MNIST addition experiments to output the probability distribution over digits, indicating the likelihood that each image in the two sequences corresponds to a specific digit (manhaeve2018deepproblog).
@@
We adopt a structured factorization of the problem as digit-wise addition with carry propagation, similar to DeepSoftLog (maene2023soft). This formulation scales linearly in the sequence length $N$. During DP, we maintain two tables: one indexed by $(i,s_{i})$ and another by $(i,c_{i})$, where $i\in\{0,\ldots,N-1\}$ denotes the digit position (with $i=0$ being the least significant digit), $s_{i}$ is the observed sum digit at position $i$, and $c_{i}$ is the carry propagated from position $i$ to $i+1$. Each entry in the DP tables represents the total probability mass of all valid latent explanations (i.e., digit assignments) leading to that state. The probability computation accounts for the distribution over the incoming carry from the previous digit, as well as the output distributions of the neural digit classifier for the two MNIST images at the current position. Since we train only on positive queries in this problem, maximizing the value function corresponds to maximizing the probability of the ground-truth sum. This probability is computed as the product of the probabilities assigned to the correct sum digit at each position.

### v2

@@ 5 Learning in DeepProofLog
Learning in DeepProofLog is equivalent to learning the optimal policy of the corresponding MDP defined in Section 4. For the empirical distribution defined by $\mathcal{D}$, the objective is defined as $J(\pi_{\lambda})=\sum_{i=1}^{N}[V^{\pi_{\lambda}}(q^{(i)},y^{(i)})]$ (Sutton and Barto 1999). Given our reward design in Equation 3 and setting the discount factor $\gamma=1$, we show in Appendix A.3 that the objective function simplifies to:
@@
Approximate Inference via Policy Gradient. When the full goal space is intractable, the mapping between SLP derivations and MDPs enables the application of reinforcement learning techniques to learn the clause selection policy. Prominent algorithms such as REINFORCE (Williams 1992), Actor-Critic (Sutton et al. 2000), and Proximal Policy Optimization (PPO) (Schulman et al. 2017) can be adapted. For example, we employ PPO to optimize the neural policy, using the standard clipped surrogate objective. Alongside the policy, we parameterize a neural value function with the same architecture to estimate returns and reduce variance during training. While value-based methods are also applicable in principle, we leave their exploration for future work.
@@ 6 Experiments
We present two sets of experiments. First, we assess our approach on a neurosymbolic task that learns perception given the background knowledge, showing that DeepProofLog (DPrL) outperforms advanced approximate inference systems in both accuracy and scalability while scaling better than exact inference systems such as DeepStochLog. The second set of experiments addresses explainable knowledge graph (KG) completion. Here, DPrL learns from data to generate high-probability proofs for positive queries and low-probability proofs for negative ones. The fact that each classification is expressed by a proof makes DPrL’s decisions fully interpretable. Comparative experiments against state-of-the-art proof-based KG systems show that our method scales to large KGs without compromising its predictive performance. All code and data will be publicly available upon publication.
@@ 6.1 MNIST Addition
NeSy Exact KLay Time T/O
@@
NeSy Approximate Scallop Acc. $92.3\pm 0.7$ T/O
@@
We compare with state-of-the-art NeSy methods on predicted sum accuracy and training time. Baselines include three exact systems: KLay (Maene et al. 2025), DeepStochLog, and DeepSoftLog (Maene and De Raedt 2023); and three approximate systems: Scallop (Huang et al. 2021), A-NeSI (van Krieken et al. 2023), and EXAL (Verreet et al. 2024). Results are summarized in Table 1, with experimental setups, hyperparameters, and additional intermediate N values provided in Appendix C.1.
The DP variant achieves accuracy comparable to existing exact NeSy systems while significantly outperforming both exact and approximate baselines in terms of training efficiency. We observed that no other previous method successfully scales to $N=100$ without timing out. In the approximate setting, our PG variant achieves the highest accuracy. It scales beyond $N=100$, demonstrating strong computational efficiency, but converges to a suboptimal policy due to the vast combinatorial search space.
6.2 Explainable Knowledge Graph Completion
{forest}
{forest}
@@ 6.2 Explainable Knowledge Graph Completion
Figure 2: Examples of multi-step proof trees discovered by DeepProofLog for the Family and WN18RR datasets.
@@
We evaluate DPrL on popular KG completion benchmarks: Family (Kok and Domingos 2007) and WN18RR (Dettmers et al. 2018). Family models familial relationships, while WN18RR is a challenging subset of WordNet with inverse relations removed (dataset statistics in Appendix C.2). All methods were tested with 200 randomly sampled negative corruptions per query.
@@
Table 2 compares DPrL to KGE baselines (ComplEx (Trouillon et al. 2016), RotatE (Sun et al. 2019)) and logic-based NeSy systems: SBR (Diligenti et al. 2017), R2N (Marra et al. 2025) and DCR (Barbiero et al. 2023). Please note that these NeSy systems would not scale to the considered datasets without relying on custom approximate grounding techniques (Ontiveros et al. 2025) We indicate these variants as $\textit{R2N}_{\text{SG}}$, $\textit{DCR}_{\text{SG}}$, and $\textit{SBR}_{\text{SG}}$. However, these fast grounding techniques require manual parameter tuning, whereas DPrL automatically learns the best exploration strategy for each dataset.
@@
DPrL uses Proximal Policy Optimization (PPO) (Schulman et al. 2017) to navigate the complex, large-scale action spaces emerging in KG reasoning. The model computes the scores $p^{\lambda}_{\text{success}}(q)$ for each query $q$ so that the standard evaluation metrics for each task (MRR, Hits@N) can be computed (see Appendix C.2 for the metric definitions). All NeSy methods, including DPrL, use RotatE output as a prior, learning a logic-based adjustment on top of it.
@@
On Family, all NeSy systems improve performance over the baseline. DPrL is particularly effective, especially for MRR and Hits@1. A similar trend is observed on WN18RR, where DPrL and SBR outperform other methods.
@@
Unlike the partially latent reasoning done by R2N or DCR, DPrL is fully interpretable. Figure 2 presents two proof trees generated by DPrL: a depth-3 proof for synset_domain_topic_of(09788237, 08441203) from WN18RR, and a proof for uncle(2266, 2252) from Family.
@@
Methods based on a complete or approximate grounding technique, like the ones proposed by (Ontiveros et al. 2025), would need to instantiate and process a combinatorially large number of ground formulas to find a proof at a high depth. In contrast, DPrL can avoid an exhaustive search by learning an efficient proof-finding policy, which navigates deep reasoning paths efficiently. For instance, on the test set of Family, the smart grounding methods of (Ontiveros et al. 2025) instantiate 18,071, 45,466, and 181,095 groundings at depths 1, 2, and 3, respectively. In contrast, DPrL needs just 12,334 ground evaluations.
@@
In conclusion, DPrL demonstrates superior scalability over other perception- and structure-learning NeSy models, such as DeepSoftLog (which fails to scale to the datasets presented here), while preserving the explainability and interpretability advantages of logic-based NeSy approaches.
@@ 7 Related Work
DeepStochLog extends SLPs by parameterizing clause probabilities via neural networks under derivation-based semantics. However, its exact inference leads to the combinatorial explosion of proof paths, making it intractable for complex tasks. It also assumes fixed output domains and conditions neural predictions on partial goals, limiting flexibility and context awareness. In contrast, DPrL factorizes the derivation distribution over resolution steps, conditions each decision on the full current goal, and supports variable output domains. This results in better scalability (linear in derivation steps and matching unification), more informed decisions, and greater expressivity.
@@
Efforts to improve the scalability of exact NeSy systems include DeepStochLog, which adopts a more efficient derivation-based semantics, and KLay (Maene et al. 2025), which introduces GPU-parallel arithmetic circuits. Approximate NeSy methods include search-based techniques (Manhaeve et al. 2021; Huang et al. 2021) and random walks with restart (Wang et al. 2013), which limit inference to parts of the proof space but risk bias. Sampling-based approaches (Verreet et al. 2025) avoid full model counting but are still limited by grounding. Variational inference (Abboud et al. 2020; Dos Martires 2021; van Krieken et al. 2023) replaces logic inference with learned predictors, trading interpretability for speed. Parametrized grounding (Castellano Ontiveros et al. 2025) offers a tunable balance between scalability and expressivity but lacks general guarantees provided by DPrL.
@@
Automated Theorem Proving (ATP) finds a proof for a query within deterministic logical frameworks. Recent advances integrate learning-based heuristics to guide proof search (Rocktäschel and Riedel 2017; Kaliszyk et al. 2018; Rawson and Reger 2019; Zombori et al. 2020; Lample et al. 2022). Unlike ATP systems, DPrL operates in a probabilistic setting and learns from labeled data. Rather than determining provability, it optimizes inference by maximizing the probability of positive queries and minimizing that of negatives.
@@
Reinforcement learning (RL) has been applied to knowledge graph (KG) completion mainly via path-finding models like DeepPath (Xiong et al. 2017) and MINERVA (Das et al. 2017), which train agents to navigate KGs by sequentially selecting entities and relations. These approaches map relation-path discovery to an MDP but are specific to KG structures. While paths can serve as explanations, they may be complex and hard to interpret. In contrast, DeepProofLog maps SLPs to MDPs, enabling RL-guided proof search across any domain expressible with SLPs, offering greater generality beyond KGs.
@@ 8 Conclusions
We introduce DeepProofLog (DPrL), a scalable and flexible NeSy framework that bridges expressive logical representations and efficient data-driven learning. DPrL leverages SLPs with derivation-based semantics and introduces a more expressive neural parameterization by conditioning on the full goal at each resolution step. A primary contribution is the novel and formal correspondence between inference and learning in our deep SLPs and policy optimization in MDPs, enabling dynamic programming and RL to tackle previously intractable NeSy problems.
The presented method shares a limitation with prior work: clause selection is restricted to the leftmost atom. While this preserves expressivity and completeness of the proofs, we plan to explore a more flexible selection, which could more efficiently prune failing derivations. Our current implementation uses standard policy gradients; future work will consider advanced RL techniques and broader evaluation settings. We also see potential in extending to possible world–based NeSy methods like Markov Logic Networks (Richardson and Domingos 2006) and in studying automatic logic program reformulations to reduce the search space (Leuschel and Bruynooghe 2002).
Acknowledgments
This work has been supported by the EU Framework Program for Research and Innovation Horizon under the Grant Agreement No 101073307 (MSCA-DN LeMuR), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101142702), and the Flemish Government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme. F. Giannini acknowledges support from the Partnership Extended PE00000013 - FAIR - Future Artificial Intelligence Research - Spoke 1 Human-centered AI and ERC-2018-ADG G.A. 834756 XAI: Science and technology for the eXplanation of AI decision making. L. De Raedt is also supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. M. Diligenti and M. Gori are also supported by the project “CONSTR: a COllectionless-based Neuro-Symbolic Theory for learning and Reasoning”, PARTENARIATO ESTESO “Future Artificial Intelligence Research - FAIR”, SPOKE 1 “Human-Centered AI” Università di Pisa, “NextGenerationEU”, CUP I53C22001380006. G. Marra acknowledges support by FWO (G033625N) and KU Leuven Research Fund (CELSA, CELSA/24/008).
Y. Jiao would like to thank Gabriele Venturato and Lennert De Smet for stimulating discussions.
@@ A.3 Alignment between MDP Objective and Cross-entropy Loss
In NeSy systems, learning is often supervised via cross-entropy between predicted success probabilities and binary labels (Winters et al. 2022). We now show that, under our MDP formulation in Section 4, the objective that maximizes the expected cumulative reward is equivalent to minimizing cross-entropy loss on a labeled dataset $\mathcal{D}=\{(q^{(i)},y^{(i)})\}_{i=1}^{N}$.
@@ C.1 MNIST Addition
We follow the same data generation process for the MNIST addition experiment as in previous works (Manhaeve et al. 2018). This involves consecutively selecting $2N$ images from the MNIST dataset and splitting them into two distinct sequences, each of length $N$. To create supervision signals, we compute the target sum values by multiplying the labels of the selected MNIST images by the corresponding powers of 10 and summing the resulting sequences. Finally, the two numbers are added together to form the final label. Each MNIST image is used only once in all sequences, resulting in $\lfloor 60000/2N\rfloor$ learning samples. They are further split in a 5:1 ratio for training and validation. The test set is generated similarly using the MNIST test partition.
@@
As in previous works, we use a standard LeNet architecture in all MNIST addition experiments to output the probability distribution over digits, indicating the likelihood that each image in the two sequences corresponds to a specific digit (Manhaeve et al. 2018).
@@
We adopt a structured factorization of the problem as digit-wise addition with carry propagation, similar to DeepSoftLog (Maene and De Raedt 2023). This formulation scales linearly in the sequence length $N$. During DP, we maintain two tables: one indexed by $(i,s_{i})$ and another by $(i,c_{i})$, where $i\in\{0,\ldots,N-1\}$ denotes the digit position (with $i=0$ being the least significant digit), $s_{i}$ is the observed sum digit at position $i$, and $c_{i}$ is the carry propagated from position $i$ to $i+1$. Each entry in the DP tables represents the total probability mass of all valid latent explanations (i.e., digit assignments) leading to that state. The probability computation accounts for the distribution over the incoming carry from the previous digit, as well as the output distributions of the neural digit classifier for the two MNIST images at the current position. Since we train only on positive queries in this problem, maximizing the value function corresponds to maximizing the probability of the ground-truth sum. This probability is computed as the product of the probabilities assigned to the correct sum digit at each position.